A professor believes hat this years english class is different compared to last years class.

Last years students got an average grade of a 78.

This years students scores were: 87, 66, 73, 72, 78, 82, 55, 92, 85, 70, 92, 75, 88, 65

which year did better?

Standard error of mean:

standard deviation:

degrees of freedom

one tail or two tailed

p value: .05

crit value:

PLEASE show all steps so I can thoroughly understand! thanks!
Answer #1

Standard error of mean:11.1069 / sqrt(14) = 2.9684


standard deviation: 11.1069


degrees of freedom = n - 1 = 14 - 1 =13


one tail or two tailed : two tailed


Test statistic,
t = (xbar - mu)/(s/sqrt(n))
t = (77.1429 - 78)/(11.1069/sqrt(14))
t = -0.289

P-value Approach
P-value = 0.7771


crit value:Critical value of t are -2.16 and 2.16.